Evaluate ∫x sinx dx
1
Expert's answer
2020-12-15T01:23:31-0500

"\int xsinxdx=\int udv\\
u=x\\ 
du=dx, \\
dv=sinxdx\\
\int dv = \int sinxdx\\
v=-cosx\\
\int udv=uv-\int vdu\\
\int xsinxdx=x(-cosx)-\int (-cosx)dx\\
\int xsinxdx=-xcosx+\int (cosx)dx\\
\int xsinxdx=-xcosx+sinx+c\\
\int xsinxdx=sinx-xcosx+c\\"